1. Find the midpoint of the segment with ends (4, -3, 1) and (-2, 5, 3).

2. Determine whether the vectors u=3i+j-2k and v=2i-4j+k are orthogonal.

3. Describe and sketch 9x^2 + 9y^2 - 4z^2=0


1
Expert's answer
2021-12-13T13:49:32-0500

1.


"x_M=\dfrac{4-2}{2}=1,"

"y_M=\dfrac{-3+5}{2}=1,"

"z_M=\dfrac{1+3}{2}=2"

"M(1,1,2)"

2.


"\vec u\cdot\vec v=3(2)+1(-4)+(-2)(1)=0"

Then the vectors "\vec u=3\vec i+\vec j-2\vec k" and "\vec v=2\vec i-4\vec j+\vec k" are orthogonal.


3.


"9x^2 + 9y^2 - 4z^2=0"

Cone

"\dfrac{z^2}{9}=\dfrac{x^2}{4}+\dfrac{y^2}{4}"

Horizontal traces are circles.

Vertical traces in the planes "x=k" and "y=k" are hyperbolas if "k\not=0" but are pairs of lines if "k=0."